OpenStudy (anonymous):

Medal!!! Solve 4((cos(3x-(pi/6))^2))-3=0, when -pi<=x<=pi.

OpenStudy (anonymous):

add 3, divide by 4, take the square root and you start with \[\cos(3x-\frac{\pi}{6})=\pm\frac{\sqrt3}{2}\]

OpenStudy (anonymous):

solve that for \(3x-\frac{\pi}{6}\) then solve the result for \(x\)